An industry is proposing to construct a new coal-fired power plant that produces 2300 MW
electricity with efficiency of 39%. Coal has a heating value of 10,750 Btu/lb, and contains a
sulfur content of 2.4%. It is expected that 95% of sulfur will be converted to SO2. The New
Source Performance Standards (NSPS) limits the SO2 emission rate from any coal-fired
power plants to 1.2 lb/ 106
Btu heat input.
a) Determine coal consumption in lb/ hr
b) Determine SO2 emission rate in lb/hr
c) Determine if the newly proposed power plant is required to install a flue gas
desulfurization (FGD) process for SO2 removal from the flue gas. What percent SO2
removal is required?
Answer #1

a) total coal power = 2300 *100/39

                                 = 5897.43 MW

        Total coal energy per hour = 21230748 *106 J

=> coal consumption in lb/ hr = 21230748 *106 /11341895

                                                = 1.872 *106 lb/hr

b)    SO2 emission rate in lb/hr = 0.024 * 1.872 *106 *0.95

                                                  = 4.268 * 104 lb/hr

c)    % SO2 removal is required = 100 - (4.268 * 104 * 1.2*10-6/ 4.268 * 104) *100

                                                      =   99.988 %